# Continuous-time q-learning in jump-diffusion models under Tsallis entropy

### v3

@@ 2.1 Exploratory formulation in reinforcement learning
    For any $k\geq 1$ and $(s,x,x^{\prime})\in[t,T]\times\mathds{R}^{2}$, there exist constant $C_{k}>0$ and $M_{k}\geq 1$ such that
    $\int_{U}|u|^{k}\pi(u|s,x)du<C_{k}(1+|x|^{M_{k}}),\quad\int_{U}l_{p}(\pi(u|s,x))\pi(u|s,x)du<C_{k}(1+|x|^{M_{k}}),$
    $\qquad\quad\int_{U}\int_{\mathds{R}}|\varphi(s,x^{\prime},u,z)||\pi(u|s,x)-\pi(u|s,x^{\prime})|^{k}du\nu(dz)\leq C_{k}|x-x^{\prime}|^{k}.$
@@
    For any $k\geq 1$, there exists a constant $C>0$ such that, for all $(t,x_{1},x_{2},u)\in[0,T]\times\mathds{R}^{2}\times U$,
@@
    and $|b(t,x,u)|+|\sigma(t,x,u)|\leq C(1+|x|)$ for all $(t,x,u)\in[0,T]\times\mathds{R}\times U$.
@@
    For any $k\geq 1$, there exists a constant $C_{k}>0$ such that, for all $(t,x_{1},x_{2},u)\in[0,T]\times\mathds{R}^{2}\times U$,
    $\int_{\mathds{R}}|\varphi(t,x_{1},u,z)-\varphi(t,x_{2},u,z)|^{k}\nu(dz)\leq C_{k}|x_{1}-x_{2}|^{k},$
    and $\int_{\mathds{R}}|\varphi(t,x,u,z)|^{k}\nu(dz)\leq C_{k}(1+|x|^{k})$ for all $(t,x,u)\in[0,T]\times\mathds{R}\times U$.
@@
    $|f(t,x,u)|+|g(x)|$ $\leq C(1+|x|^{k}+|u|^{k}),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall(t,x,u)\in[0,T]\times\mathds{R}\times U.$
@@ 2.2 Weak convergence of sampled dynamics to relaxed version
This subsection shows that when actions are sampled at a discrete time grid, the resulting state process (2.1) converges weakly to the relaxed version of state process (2.1) as the time stepsize goes to zero.
The following lemma establishes the existence and uniqueness of strong solution to the sampled dynamics (2.1) and the relaxed version of dynamics (2.1).
@@
Let Assumptions (A1) and (A2) hold. Consider $(t,x)\in[0,T]\times\mathds{R}$, $\pi\in\Pi_{t}$ and $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime},\mathbb{Q})$ being the probability space specified in (2.3). Then, we have
@@
    for any grid $G_{t:T}$, the SDE (2.1) with $X_{t}^{\pi}=x$ has a unique strong solution. Moreover, for all $k\geq 2$, there exists a constant $C_{k}>0$ such that, for any grid $G_{t:T}$ and $s\in[t,T]$,
    $\mathbb{E}^{\mathbb{Q}}\left[\left|X^{\pi}_{s}\right|^{k}\right]\leq C_{k}(1+x^{k})e^{C_{k}s}.$
@@
For simplicity, let $C>0$ be a generic constant independent of $(s,x,u)$, and $C_{k}$ be a generic constant independent of $(s,x,u)$ but depending on $k$, which may be different from line to line. We first show item (i). Fix a grid $G_{t:T}$. Note that $x\in\mathds{R}$ is deterministic, it follows from Assumptions (A1) and (A2), together with Theorem 3.1 and Theorem 3.2 in Kunita (2004) that, the SDE (2.1) with $X_{t}^{\pi}=x$ has a unique strong solution on $s_{0},s_{1}]$ satisfying
$\sup_{s\in[s_{0},s_{1}]}\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s}^{\pi}\right|^{k}\right]\leq C_{k}(1+x^{k}),\quad\forall k\geq 2.$
We next assume that the SDE (2.1) with $X_{t}^{\pi}=x$ has a unique strong solution on $s_{0},s_{i}]$ for some $i=1,\ldots,n-1$ satisfying $\sup_{s\in[s_{0},s_{i}]}\mathbb{E}^{\mathbb{Q}}[|X_{s}^{\pi}|^{k}]\leq C_{k}(1+x^{k})$. On the time interval $s_{i},s_{i+1}]$, as $\mathbb{E}^{\mathbb{Q}}[|X_{s_{i}}^{\pi}|^{k}]\leq C_{k}(1+x^{k})<\infty$, by applying Theorem 3.1 and Theorem 3.2 in Kunita (2004) again, we get the existence and uniqueness of strong solution of SDE (2.1). Furthermore, it holds that
$\sup_{s\in[s_{i},s_{i+1}]}\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s}^{\pi}\right|^{k}\right]\leq C_{k}\left\{1+\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s_{i}}^{\pi}\right|^{k}\right]\right\}\leq C_{k}(1+x^{k}),\quad\forall k\geq 2.$ (2.10)
By induction, we establish the well-posedness of SDE (2.1) in the sense of strong solution.
Next, we handle item (ii). From the proof of Lemma 2 in Jia and Zhou (2022b), for all $(s,x_{1},x_{2})\in[t,T]\times\mathds{R}^{2}$, it follows that
$|\tilde{b}(s,x_{1},\pi)-\tilde{b}(s,x_{2},\pi)|+|\tilde{\sigma}(s,x_{1},\pi)-\tilde{\sigma}(s,x_{2},\pi)|$ $\leq C|x_{1}-x_{2}|,$ (2.11)
and for all $(s,x)\in[t,T]\times\mathds{R}$,
$|\tilde{b}(s,x,\pi)|+|\tilde{\sigma}(s,x,\pi)|$ $\leq C(1+|x|).$ (2.12)
For $({t},x,x^{\prime})\in[0,T]\times\mathds{R}^{2}$, it follows from $\pi\in\Pi_{t}$ and Assumption (A2) that
$\left|\int_{U}\int_{\mathds{R}}\varphi(t,x,u,z)\nu(dz)\pi(u|t,x)du-\int_{U}\int_{\mathds{R}}\varphi(t,x^{\prime},u,z)\nu(dz)\pi(u|t,x^{\prime})du\right|$
$\qquad\leq\int_{U}\int_{\mathds{R}}\left|\varphi(t,x,u,z)-\varphi(t,x^{\prime},u,z)\right|\nu(dz)\pi(u|t,x)du$
$\qquad\quad+\int_{U}\int_{\mathds{R}}\varphi(t,x^{\prime},u,z)\nu(dz)\left|\pi(u|t,x)-\pi(u|t,x^{\prime})\right|du$
@@
and for $(t,x)\in[0,T]\times\mathds{R}$,
$\int_{U}\int_{\mathds{R}}\left|\varphi(t,x,u,z)\right|^{k}\nu(dz)\pi(u|t,x)du\leq C_{k}(1+|x|^{k}).$ (2.14)
In view of (2.11)-(2.14) and Theorem 3.1 in Kunita (2004), the SDE (2.1) with $\tilde{X}_{t}^{\pi}=x$ admits a unique strong solution. ∎
Similar to Jia et al. (2025), we introduce the following notations for the convergence analysis. For each $m\in\mathbb{N}$ and $k>0$, a function $v:[0,T]\times\mathds{R}\to\mathds{R}$ belongs to $C^{m}_{k}([0,T]\times\mathds{R})$ provided for all $r,\ell\in\mathbb{N}_{0}$ satisfying $2r+\ell\leq m$, the partial derivative $\partial_{t}^{r}\partial_{x}^{\ell}v$ exists, is continuous on $0,T]\times\mathds{R}$, and all such derivatives satisfy a polynomial growth condition in $x$:
$\|v\|_{C_{k}^{m}}:=\sum_{\begin{subarray}{c}2r+\ell\leq m\end{subarray}}\sup_{(t,x)\in[0,T]\times\mathds{R}}\frac{\left|\partial_{t}^{r}\partial_{x}^{\ell}v(t,x)\right|}{1+|x|^{k}}<\infty.$
Let $C^{m,0}_{k}([0,T]\times\mathds{R}\times U)$ be the space of functions $v:[0,T]\times\mathds{R}\times U\to\mathds{R}$ such that, for all $r,s\in\mathbb{N}_{0}$ satisfying $2r+s\leq m$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}v(t,x,u)$ exists, is continuous in $(t,x)$ for all $(t,x,u)\in[0,T]\times\mathds{R}\times U$, and all such derivatives satisfy a polynomial growth condition in $x$ and $u$:
$\|v\|_{C_{k}^{m,0}}:=\sum_{\begin{subarray}{c}2r+s\leq m\end{subarray}}\sup_{(t,x,u)\in[0,T]\times\mathds{R}\times U}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}v(t,x,u)\right|}{1+|x|^{k}+|u|^{k}}<\infty.$
@@
    $b,\sigma^{2}\in C^{2,0}_{k}([0,T]\times\mathds{R}\times U)$, $\varphi(\cdot,\cdot,u,z)\in C^{1,2}([0,T]\times\mathds{R})$ for any $(u,z)\in U\times\mathds{R}$, and there exists some constant $C>0$ such that, for all $(t,x,u,z)\in[0,T]\times\mathds{R}\times U\times\mathds{R}$,
    $|\varphi_{t}(t,x,u,z)|+|\varphi_{x}(t,x,u,z)|+|\varphi_{xx}(t,x,u,z)|\leq C.$
@@
    for $\pi\in\Pi_{0}$, $\tilde{b},\tilde{\sigma},\varphi,\nu$ are sufficiently regular in the sense that, for any $h\in C_{k}^{4}(\mathds{R})$ and $t^{\prime}\in[0,T]$, the following PDE, for $(t,x)\in[0,t^{\prime})\times\mathds{R}$,
    $\begin{cases}\displaystyle v_{t}(t,x)+\tilde{b}(t,x,\pi)v_{x}(t,x)+\frac{\tilde{\sigma}^{2}(t,x,\pi)}{2}v_{xx}(t,x)\\ \displaystyle\quad+\int_{U}\int_{\mathds{R}}\left(v\left(t,x+\varphi(t,x,u,z)\right)-v(t,x)\right)\nu(dz)\pi(u|t,x)du=0,\\[6.00006pt] \displaystyle v(t^{\prime},x)=h(x),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x\in\mathds{R},\end{cases}$ (2.15)
    has a unique solution $v^{h}(\cdot,\cdot;t^{\prime})\in C_{k}^{4}([0,t^{\prime}]\times\mathds{R})$. Moreover, there exists a constant $C>0$ depending only on $(T,k,\tilde{b},\tilde{\sigma},\varphi,\pi)$ such that $\|v^{h}(\cdot,\cdot;t^{\prime})\|_{C_{k}^{4}}\leq C\|h\|_{C_{k}^{4}}$.
@@
    $f\in C_{k}^{2,0}$, $\tilde{f}\in C_{k}^{2}$ and $\tilde{f}(t,\cdot),g\in C_{k}^{4}(\mathds{R})$ for any $t\in[0,T]$.
@@
Let Assumptions (A1), (A2), (A4) and (A5) hold with $k\geq 2$. For $\pi\in\Pi_{0}$, consider the sampled dynamics (2.1) and the relaxed version of dynamics (2.1) with initial value $x\in\mathds{R}$. Then, for any $h\in C_{k}^{4}(\mathds{R})$, there exists a constant $C_{k}>0$ such that, for all grids $G_{0:T}$,
@@
Here, $\|G_{0:T}\|:=\max_{0\leq i\leq n-1}(s_{i+1}-s_{i})$ is the mesh size of the time grid $G_{0:T}=\{0=s_{0}<s_{1}<\cdots<s_{n}=T\}$.
@@
By Assumption (A5), it holds that, for $h\in C_{k}^{4}(\mathds{R})$ and $t^{\prime}=s_{j}$ with $1\leq j\leq n$, the PDE (2.15) has a unique solution $v^{h}(\cdot,\cdot;s_{j})\in C_{k}^{4}([0,t^{\prime}]\times\mathds{R})$. For simplicity, we omit $h$, $s_{j}$, and $\pi$. It follows from the Feymann-Kac’s formula that $\mathbb{E}^{\mathbb{Q}}[h(\tilde{X}_{t}^{\pi})]=h(0,x)$. Then, we have
$\mathbb{E}^{\mathbb{Q}}\left[h(X_{s_{j}})-h(\tilde{X}_{s_{j}})\right]$ $=\mathbb{E}^{\mathbb{Q}}\left[v(s_{j},X_{s_{j}})-v(0,x)\right]$
$=\mathbb{E}^{\mathbb{Q}}\left[\sum_{i=1}^{j}\left(v(s_{i},X_{s_{i}})-v(s_{i-1},X_{s_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}^{\mathbb{Q}}\left[I_{i}\right].$ (2.17)

### v4

@@ 2.1 Exploratory formulation in reinforcement learning
    For any $k\geq 1$ and $(s,x,x^{\prime})\in[t,T]\times\mathds{R}^{2n}$, there exist constants $C,C_{k},C_{p}>0$ and $M_{p}\geq 1$ such that
    $\int_{U}|u|^{k}\pi(u|s,x)du<C_{k}(1+|x|^{k}),\quad\int_{U}l_{p}(\pi(u|s,x))\pi(u|s,x)du<C_{p}(1+|x|^{M_{p}}),$
    $\qquad\quad\int_{U}\int_{\mathds{R}^{n}}|\varphi(s,x^{\prime},u,z)\lambda(s,x^{\prime},u)||\pi(u|s,x)-\pi(u|s,x^{\prime})|\nu(dz)du\leq C|x-x^{\prime}|.$
@@
    There exists $C>0$ such that, for all $(t,x_{1},x_{2},x,u)\in[0,T]\times\mathds{R}^{3n}\times U$,
@@
    $\left|b(t,x,u)\right|+\left|\sigma(t,x,u)\right|\leq C(1+|x|+|u|)$ .
@@
    For any $k\geq 1$, there exist $C,C_{k}>0$ such that, for all $(t,x_{1},x_{2},u)\in[0,T]\times\mathds{R}^{2n}\times U$,
    $\int_{\mathds{R}^{n}}|\varphi(t,x_{1},u,z)-\varphi(t,x_{2},u,z)||\lambda(t,x_{1},u)|\nu(dz)$ $\leq C|x_{1}-x_{2}|,$
    $\int_{\mathds{R}^{n}}|\varphi(t,x_{2},u,z)||\lambda(t,x_{1},u)-\lambda(t,x_{2},u)|\nu(dz)$ $\leq C|x_{1}-x_{2}|,$
    $\int_{\mathds{R}}|\varphi(t,x,u,z)|^{k}\lambda(t,x,u)\nu(dz)$ $\leq C_{k}(1+|x|^{k}+|u|^{k}).$
@@
    $|f(t,x,u)|+|g(x)|$ $\leq C(1+|x|^{k}+|u|^{k}),\penalty 10000\ \penalty 10000\ \forall(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U.$
@@ 2.2 Weak convergence of sampled dynamics to relaxed version
This subsection shows that when actions are sampled at a discrete time grid, the resulting state process (2.1) converges weakly to the relaxed version of state process (2.1) as the time stepsize goes to zero. The following lemma establishes the existence and uniqueness of strong solution to the sampled dynamics (2.1) and the relaxed version of dynamics (2.1).
@@
Let Assumptions (A1) and (A2) hold. Consider $(t,x)\in[0,T]\times\mathds{R}^{n}$, $\pi\in\Pi_{t}$ and $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime},\mathbb{Q})$ which is the probability space specified in (2.3). Then, we have
@@
    for any grid $G_{t:T}$, the SDE (2.1) with $X_{t}^{\pi}=x$ has a unique strong solution. Moreover, for all $k\geq 2$, there exists a constant $C_{k}>0$ such that $\mathbb{E}^{\mathbb{Q}}[|X^{\pi}_{s}|^{k}]\leq C_{k}(1+|x|^{k})e^{C_{k}s}$ for any grid $G_{t:T}$ and $s\in[t,T]$.
@@
For simplicity, let $C>0$ be a generic constant independent of $(s,x,u)$, and $C_{k}$ be a generic constant independent of $(s,x,u)$ but depending on $k$, which may differ from line to line. We first show item (i). Fix a grid $G_{t:T}$. By using Assumption (A2), we deduce that, for any $(s,x_{1},x_{2},u)\in[t,T]\times\mathds{R}^{2n}\times U$,
$\left|\int_{\mathds{R}^{n}}\left(\varphi(s,x_{1},u,z)\lambda(s,x_{1},u)-\varphi(s,x_{2},u,z)\lambda(s,x_{2},u)\right)\nu(dz)\right|$
$\leq\int_{\mathds{R}^{n}}|\varphi(s,x_{1},u,z)-\varphi(s,x_{2},u,z)|\lambda(s,x_{1},u)\nu(dz)$
$\quad+\int_{\mathds{R}^{n}}|\varphi(s,x_{2},u,z)||\lambda(s,x_{1},u)-\lambda(s,x_{2},u)|\nu(dz)\leq C|x_{1}-x_{2}|.$ (2.10)
As $x\in\mathbb{R}$ is deterministic, in view of (2.2) and Assumptions (A1)–(A2) and following arguments in the proof of Theorem 1.2 in Graham (1992) and Theorem 3.2 in Kunita (2004), it holds that SDE (2.1) with $X_{t}^{\pi}=x$ admits a unique strong solution on $s_{0},s_{1}]$ satisfying
$\sup_{s\in[s_{0},s_{1}]}\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s}^{\pi}\right|^{k}\right]\leq C_{k}(1+|x|^{k}),\quad\forall k\geq 2.$
We next assume that SDE (2.1) with $X_{t}^{\pi}=x$ has a unique strong solution on $s_{0},s_{i}]$ for some $i=1,\ldots,n_{0}-1$ satisfying $\sup_{s\in[s_{0},s_{i}]}\mathbb{E}^{\mathbb{Q}}[|X_{s}^{\pi}|^{k}]\leq C_{k}(1+|x|^{k})$. On the time interval $s_{i},s_{i+1}]$, as $\mathbb{E}^{\mathbb{Q}}[|X_{s_{i}}^{\pi}|^{k}]\leq C_{k}(1+|x|^{k})<\infty$, we have, for all $s\in[s_{i},s_{i+1}]$,
$\mathbb{E}^{\mathbb{Q}}\left[\left|u_{s}^{\pi}\right|^{k}\right]=\mathbb{E}^{\mathbb{Q}}\left[\int_{U}|u|^{k}\pi(u|s_{i},X_{s_{i}}^{\pi})du\right]$ $\leq C_{k}\left\{1+\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s_{i}}^{\pi}\right|^{k}\right]\right\}\leq C_{k}(1+|x|^{k}).$ (2.11)
By applying (2.11) and the same arguments in Graham (1992) and Kunita (2004) again, we get the existence and uniqueness of strong solution to SDE (2.1). Furthermore, it holds that
$\sup_{s\in[s_{i},s_{i+1}]}\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s}^{\pi}\right|^{k}\right]\leq C_{k}\left\{1+\mathbb{E}^{\mathbb{Q}}\left[\left|X_{s_{i}}^{\pi}\right|^{k}\right]\right\}\leq C_{k}(1+|x|^{k}),\quad\forall k\geq 2.$
By induction, we establish the well-posedness of SDE (2.1) with the strong solution.
Next, we handle item (ii). From the proof of Lemma 2 in Jia and Zhou (2022b), for all $(s,x_{1},x_{2})\in[t,T]\times\mathds{R}^{2n}$, it follows that
$|\tilde{b}(s,x_{1};\pi)-\tilde{b}(s,x_{2};\pi)|+|\tilde{\sigma}(s,x_{1};\pi)-\tilde{\sigma}(s,x_{2};\pi)|$ $\leq C|x_{1}-x_{2}|,$ (2.12)
and for all $(s,x)\in[t,T]\times\mathds{R}^{n}$,
$|\tilde{b}(s,x;\pi)|+|\tilde{\sigma}(s,x;\pi)|$ $\leq C(1+|x|).$ (2.13)
For $({t},x,x^{\prime})\in[0,T]\times\mathds{R}^{2n}$, it follows from $\pi\in\Pi_{t}$, Assumption (A2) and (2.2) that
$\left|\int_{U}\int_{\mathds{R}}\varphi(t,x,u,z)\lambda(t,x,u)\nu(dz)\pi(u|t,x)du-\int_{U}\int_{\mathds{R}}\varphi(t,x^{\prime},u,z)\lambda(t,x^{\prime},u)\nu(dz)\pi(u|t,x^{\prime})du\right|$
$\qquad\leq\int_{U}\int_{\mathds{R}}\left|\varphi(t,x,u,z)\lambda(t,x,u)-\varphi(t,x^{\prime},u,z)\lambda(t,x^{\prime},u)\right|\nu(dz)\pi(u|t,x)du$
$\qquad\quad+\int_{U}\int_{\mathds{R}}|\varphi(t,x^{\prime},u,z)\lambda(t,x^{\prime},u)\left|\pi(u|t,x)-\pi(u|t,x^{\prime})\right|\nu(dz)du$
@@
and for $(t,x)\in[0,T]\times\mathds{R}^{n}$,
$\int_{U}\int_{\mathds{R}^{n}}|\varphi(t,x,u,z)|^{k}\lambda(t,x,u)\nu(dz)\pi(u|t,x)du$ $\leq\int_{U}C_{k}(1+|x|^{k}+|u|^{k})\pi(u|t,x)du$
$\leq C_{k}(1+|x|^{k}).$ (2.15)
In view of (2.12)-(2.2) and Theorem 1.2 in Graham (1992), the SDE (2.1) with $\tilde{X}_{t}^{\pi}=x$ admits a unique strong solution. ∎
Similar to Jia et al. (2025), we introduce the following notations for the convergence analysis. For each $\ell\in\mathbb{N}$ and $k>0$, a function $v:[0,T]\times\mathds{R}^{n}\to\mathds{R}$ belongs to $C^{\ell}_{k}([0,T]\times\mathds{R}^{n})$ provided for all $\ell_{1},\ell_{2}\in\mathbb{N}_{0}$ satisfying $2\ell_{1}+\ell_{2}\leq\ell$, the partial derivative $\partial_{t}^{\ell_{1}}\partial_{x}^{\ell_{2}}v$ exists, is continuous on $0,T]\times\mathds{R}^{n}$, and all such derivatives satisfy a polynomial growth condition in $x$:
$\|v\|_{C_{k}^{\ell}}:=\sum_{\begin{subarray}{c}2\ell_{1}+\ell_{2}\leq\ell\end{subarray}}\sup_{(t,x)\in[0,T]\times\mathds{R}^{n}}\frac{\left|\partial_{t}^{\ell_{1}}\partial_{x}^{\ell_{2}}v(t,x)\right|}{1+|x|^{k}}<\infty.$
Let $C^{\ell,0}_{k}([0,T]\times\mathds{R}^{n}\times U)$ be the space of functions $v:[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}$ such that, for all $\ell_{1},\ell_{2}\in\mathbb{N}_{0}$ satisfying $2\ell_{1}+\ell_{2}\leq\ell$, the partial derivative $\partial_{t}^{\ell_{1}}\partial_{x}^{\ell_{2}}v(t,x,u)$ exists, is continuous in $(t,x)$ for all $(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U$, and all such derivatives satisfy a polynomial growth condition in $x$ and $u$:
$\|v\|_{C_{k}^{\ell,0}}:=\sum_{\begin{subarray}{c}2\ell_{1}+\ell_{2}\leq\ell\end{subarray}}\sup_{(t,x,u)\in[0,T]\times\mathds{R}^{n}\times U}\frac{\left|\partial_{t}^{\ell_{1}}\partial_{x}^{\ell_{2}}v(t,x,u)\right|}{1+|x|^{k}+|u|^{k}}<\infty.$
@@
    $b,\sigma^{2}\in C^{2,0}_{k}([0,T]\times\mathds{R}^{n}\times U)$, $\lambda(\cdot,\cdot,u),\varphi(\cdot,\cdot,u,z)\in C^{1,2}([0,T]\times\mathds{R}^{n})$ for any $(u,z)\in U\times\mathds{R}^{n}$, and there exists some constant $C>0$ such that, for all $(t,x,u,z)\in[0,T]\times\mathds{R}^{n}\times U\times\mathds{R}^{n}$,
    $|\varphi_{t}(t,x,u,z)|+|\varphi_{x}(t,x,u,z)|+|\varphi_{xx}(t,x,u,z)|\leq C,$
    $|\lambda(t,x,u)|+|\lambda_{t}(t,x,u)|+|\lambda_{x}(t,x,u)|+|\lambda_{xx}(t,x,u)|\leq C.$
@@
    for $\pi\in\Pi_{0}$, $\tilde{b},\tilde{\sigma},\varphi,\nu$ are sufficiently regular in the sense that, for any $h\in C_{k}^{4}(\mathds{R}^{n})$ and $t^{\prime}\in[0,T]$, the following PDE, for $(t,x)\in[0,t^{\prime})\times\mathds{R}^{n}$,
    $\begin{cases}\displaystyle v_{t}(t,x)+\tilde{b}(t,x;\pi)\cdot v_{x}(t,x)+\frac{1}{2}\tilde{\sigma}^{2}(t,x;\pi)\circ v_{xx}(t,x)\\ \displaystyle\quad+\int_{U}\int_{\mathds{R}^{n}}\left(v\left(t,x+\varphi(t,x,u,z)\right)-v(t,x)\right)\lambda(t,x,u)\nu(dz)\pi(u|t,x)du=0,\\[6.00006pt] \displaystyle v(t^{\prime},x)=h(x),\penalty 10000\ \penalty 10000\ \forall x\in\mathds{R}^{n},\end{cases}$ (2.16)
    has a unique solution $v^{h}(\cdot,\cdot;t^{\prime})\in C_{k}^{4}([0,t^{\prime}]\times\mathds{R}^{n})$. Moreover, there exists a constant $C>0$ depending only on $(T,k,\tilde{b},\tilde{\sigma},\varphi,\pi)$ such that $\|v^{h}(\cdot,\cdot;t^{\prime})\|_{C_{k}^{4}}\leq C\|h\|_{C_{k}^{4}}$.
@@
    $f\in C_{k}^{2,0}$, $\tilde{f}\in C_{k}^{2}$ and $\tilde{f}(t,\cdot),g\in C_{k}^{4}(\mathds{R}^{n})$ for any $t\in[0,T]$.
Remark 2.2.
Note that Assumption (A5) involves the solvability of PDE (2.16), which must be checked in each specific case. Here, we can provide some examples in which Assumption (A5) holds. Let us consider the case where the jump size function $\varphi(t,x,u,z)$ is independent of the control variable $u$ (thus we omit $u$ in the functions $\varphi$ and $\lambda$) and satisfies $\varphi(t,x,0)=0$ for all $(t,x)\in[0,T]\times\mathds{R}^{n}$, and the function $\lambda(t,x)\equiv 1$. Furthermore, for $\pi\in\Pi_{0}$, assume that the functions $\tilde{b},\tilde{\sigma}\in C^{4}([0,T]\times\mathds{R}^{n}),\varphi\in C^{4}([0,T]\times\mathds{R}^{n}\times\mathds{R}^{n})$ are all bounded and have bounded partial derivatives. Then, using the stochastic flow analysis (c.f. Theorem 3.4.1 and Theorem 3.4.2 in Kunita 2019), we can verify that, for any $h\in C_{k}^{4}(\mathds{R}^{n})$ and $t^{\prime}\in[0,T]$, the function $v^{h}(t,x):=\mathbb{E}[h(\tilde{X}_{t^{\prime}}^{\pi})])\in C_{k}^{4}([0,t^{\prime}]\times\mathds{R}^{n})$ is the unique solution to PDE (2.16), where $\tilde{X}^{\pi}$ is given by (2.1) with $\tilde{X}_{t}^{\pi}=x$. Moreover, there exists a constant $C>0$ such that $\|v^{h}\|_{C_{k}^{4}}\leq C\|h\|_{C_{k}^{4}}$.
@@
Let Assumptions (A1), (A2), (A4) and (A5) hold with $k\geq 2$. For $\pi\in\Pi_{0}$, consider the sampled dynamics (2.1) and the relaxed version of dynamics (2.1) with initial value $x\in\mathds{R}^{n}$. Then, for any $h\in C_{k}^{4}(\mathds{R}^{n})$, there exists a constant $C_{k}>0$ such that, for all grids $G_{0:T}$,
@@
Here, $\|G_{0:T}\|:=\max_{0\leq i\leq n_{0}-1}(s_{i+1}-s_{i})$ is the mesh size of the time grid $G_{0:T}=\{0=s_{0}<s_{1}<\cdots<s_{n_{0}}=T\}$.
@@
Without loss of generality, we only consider $n=1$ for notational simplicity. By Assumption (A5), for $h\in C_{k}^{4}(\mathds{R}^{n})$ and $t^{\prime}=s_{j}$ with $1\leq j\leq n_{0}$, Eq. (2.16) has a unique solution $v^{h}(\cdot,\cdot;s_{j})\in C_{k}^{4}([0,t^{\prime}]\times\mathds{R}^{n})$. For simplicity, we omit $h$, $s_{j}$, and $\pi$. It follows from the Feymann-Kac’s formula that $\mathbb{E}^{\mathbb{Q}}[h(\tilde{X}_{t})]=v(0,x)$. Then, we have
$\mathbb{E}^{\mathbb{Q}}\left[h(X_{s_{j}})-h(\tilde{X}_{s_{j}})\right]$ $=\mathbb{E}^{\mathbb{Q}}\left[v(s_{j},X_{s_{j}})-v(0,x)\right]=\mathbb{E}^{\mathbb{Q}}\left[\sum_{i=1}^{j}\left(v(s_{i},X_{s_{i}})-v(s_{i-1},X_{s_{i-1}})\right)\right]$
$=:\sum_{i=1}^{j}\mathbb{E}^{\mathbb{Q}}\left[I_{i}\right].$ (2.18)